# LMPOcc: 3D Semantic Occupancy Prediction Utilizing Long-Term Memory Prior from Historical Traversals

### v2

@@ IV-A Datasets and Metrics
We evaluate our approach on the Occ3D-nuScenes benchmark [25], which extends the widely adopted large-scale autonomous driving dataset nuScenes [34]. This benchmark comprises 700 training scenes and 150 validation scenes, each with 40 annotated samples captured at 2Hz over 20-second sequences. The dataset spans a spatial domain of $\text{X}\in[-40m,40m]$ and $\text{Y}\in[-40m,40m]$ for horizontal dimensions, with vertical coverage from $Z\in[-1m,5.4m]$. The occupancy annotations in this dataset are represented as axis-aligned voxels with 0.4m edge length, achieving a resolution of $200\times 200\times 16$ voxels. Each voxel is labeled with one of 17 semantic categories or marked as free space (non-occupied). In our work, $Dynamic$ represents dynamic semantic categories, which encompass $others$, $barrier$, $bicycle$, $bus$, $car$, $construction\ vehicle$, $motorcycle$, $pedestrian$, $traffic\ cone$, $trailer$ and $truck$. $Static$ denotes static semantic categories, which comprise $driveable\ surface$, $other\ flat$, $sidewalk$, $terrain$, $manmade$ and $vegetation$. For performance evaluation, we employ the mean Intersection-Over-Union (mIoU) metric aggregated across all semantic classes.
@@ IV-B Implementation Details
We employ FlashOcc [24] and DHD [28] as baseline models. During training, we initialize the models with their pretrained weights and freeze the parameters preceding the current latent feature. After integrating our LMOP module, we train for another 24 epochs while maintaining identical experimental configurations to the baseline setup. We disable BEV-space data augmentation to prevent misalignment between current features and prior features. The channel dimension of the global occupancy in LMPOcc is computed as the height of the occupancy multiplied by the number of semantic categories, specifically $16\times 18$, while other configurations regarding the global map remain consistent with those in Neural Map Prior [1]. All models are trained with a batch size of 4 on 6 NVIDIA A100 GPUs.
@@ IV-C Main Results
Figure 3: Visualization results of LMPOcc and DHD. The left, center, and right columns demonstrate the performance under sunny, rainy, and night-time conditions respectively. LMPOcc demonstrates significant performance enhancement over the baseline across three distinct operational scenarios. The color-semantic category correspondence is detailed in Table I.
@@ IV-D Ablation Studies
Discussion of Dynamic Targets. Given the temporal inconsistencies of dynamic targets at the same geographic location across distinct observation periods, we conduct analysis on the handling of dynamic targets within the global prior in our framework. We propose two distinct methodologies for eliminating dynamic elements within the prior framework. After applying the argmax operation to the occupancy logits, the semantic label for each voxel can be obtained. We create free masks and dynamic masks to, respectively, mask out the free components and dynamic components from the occupancy logits. The dynamic components comprise dynamic semantic categories, as detailed in Sec. IV-A. The first method sets the dynamic components in the occupancy logits as zero, denoted as $Removing\ Dynamic\ v1$ in Table VI. The second method randomly selects the free components in the occupancy logits and replaces the dynamic components with them, denoted as $Removing\ Dynamic\ v2$ in Table VI. Both $Removing\ Dynamic\ v1$ and $Removing\ Dynamic\ v2$ achieve the effect demonstrated in Fig. 6(b).
@@
Figure 6: Visualization results of Retaining Dynamic and Retaining Dynamic. Figure (a) demonstrates the effects when retaining dynamic elements, while both Removing Dynamic v1 and Removing Dynamic v2 methodologies achieve the visual results shown in (b). The color-semantic category correspondence is detailed in Table I.
@@ IV-E Visualization
The visualization results of LMPOcc-S are presented in Fig. 3. We demonstrate three operational scenarios: sunny, rainy, and night-time conditions. Across all scenarios, LMPOcc consistently outperforms the baseline, particularly in static object categories such as manmade and vegetation. The results in the middle column demonstrate that LMPOcc maintains robust detection of static targets even under severe image degradation conditions, which is attributed to its effective utilization of long-term memory prior. Our global occupancy is shown in Fig. 4. While enhancing local perception capabilities, LMPOcc can construct occupancy representations for large-scale scenes. As shown in Fig. 5, three collaborative agents construct the global occupancy map through crowdsourcing.
@@ V CONCLUSION
We propose a novel 3D occupancy prediction framework LMPOcc, which effectively leverages long-term memory priors from historical traversals to enhance local perception performance while establishing a unified global occupancy representation. It thereby enables city-level occupancy construction through crowdsourcing. Concretely, we introduce a model-agnostic prior format to ensure compatibility across different occupancy prediction baselines, coupled with an efficient lightweight Current-Prior Fusion module to adaptively integrate prior features and current features. LMPOcc achieves state-of-the-art performance in local occupancy prediction, with experimental validation on the Occ3D-nuScenes benchmark. Future research could conduct in-depth investigations into dynamic object processing within prior information and develop learnable optimization frameworks for global occupancy.

### v3

@@ IV-A Datasets and Metrics
We evaluate our approach on the Occ3D-nuScenes benchmark [19], which extends the widely adopted large-scale autonomous driving dataset nuScenes [1]. This benchmark comprises 700 training scenes and 150 validation scenes, each with 40 annotated samples captured at 2Hz over 20-second sequences. The dataset spans a spatial domain of $\text{X}\in[-40m,40m]$ and $\text{Y}\in[-40m,40m]$ for horizontal dimensions, with vertical coverage from $Z\in[-1m,5.4m]$. The occupancy annotations in this dataset are represented as axis-aligned voxels with 0.4m edge length, achieving a resolution of $200\times 200\times 16$ voxels. Each voxel is labeled with one of 17 semantic categories or marked as free space (non-occupied). In our work, $Dynamic$ represents dynamic semantic categories, which encompass othersothers, barrierbarrier, bicyclebicycle, busbus, carcar, constructionvehicleconstruction\ vehicle, motorcyclemotorcycle, pedestrianpedestrian, trafficconetraffic\ cone, trailertrailer and trucktruck. StaticStatic denotes static semantic categories, which comprise driveablesurfacedriveable\ surface, otherflatother\ flat, sidewalksidewalk, terrainterrain, manmademanmade and vegetationvegetation. For performance evaluation, we employ the mean Intersection-Over-Union (mIoU) metric aggregated across all semantic classes.
@@ IV-B Implementation Details
We employ FlashOcc [25] and DHD [22] as baseline models. During training, we initialize the models with their pretrained weights and freeze the parameters preceding the current latent feature. After integrating our LMOP module, we train for another 24 epochs while maintaining identical experimental configurations to the baseline setup. We disable BEV-space data augmentation to prevent misalignment between current features and prior features. The channel dimension of the global occupancy in LMPOcc is computed as the height of the occupancy multiplied by the number of semantic categories, specifically $16\times 18$, while other configurations regarding the global map remain consistent with those in Neural Map Prior [23]. All models are trained with a batch size of 4 on 6 NVIDIA A100 GPUs. The 3D open vocabulary maps are constructed using 6-view images per occupancy frame, with depth sampled at intervals $\Delta d=0.1\,m$ up to a maximum depth $D_{\max}=100\,m$.
@@ IV-D Ablation Studies
Figure 6: Visualization results of LMPOcc and DHD. It is a low-visibility rainy scene. LMPOcc leverages long-term memory priors to detect objects not visible in current sensory observations, demonstrating significant improvement over the baseline. The color-semantic category correspondence is detailed in Table I.
Figure 7: Visualization results of the 3D open vocabulary map. Depth information is derived from the occupancy map, and the outdoor open-vocabulary map is constructed based on the OpenGraph method. The bottom row shows scene-level occupancy grounding for specific queries (e.g., “parked trucks”).
Discussion of Dynamic Targets. Given the temporal inconsistencies of dynamic targets at the same geographic location across distinct observation periods, we conduct analysis on the handling of dynamic targets within the global prior in our framework. We propose two distinct methodologies for eliminating dynamic elements within the prior framework. After applying the argmax operation to the occupancy logits, the semantic label for each voxel can be obtained. We create free masks and dynamic masks to, respectively, mask out the free components and dynamic components from the occupancy logits. The dynamic components comprise dynamic semantic categories, as detailed in Sec. IV-A. The first method sets the dynamic components in the occupancy logits as zero, denoted as RemovingDynamicv1Removing\ Dynamic\ v1 in Table VI. The second method randomly selects the free components in the occupancy logits and replaces the dynamic components with them, denoted as RemovingDynamicv2Removing\ Dynamic\ v2 in Table VI.
@@ IV-E Visualization
Our global occupancy is shown in Fig. 4. While enhancing local perception capabilities, LMPOcc can construct global occupancy for large-scale scenes. As shown in Fig. 5, three collaborative agents construct the global occupancy map through crowdsourcing. The visualization results of LMPOcc-S are presented in Fig. 6. It shows a rainy, low-visibility scene where LMPOcc leverages long-term memory priors to detect objects missing in current views, significantly outperforming the baseline.
IV-F Construction of 3D Open Vocabulary Maps
In Fig. 7, we showcase the results of a 3D open-vocabulary map constructed with OpenGraph [3], utilizing depth generated from the occupancy map for a sequence of multi-view images. The first row shows the image sequence used for building the 3D map, the second row illustrates the depth extracted from the occupancy map for each image via ray casting, and the third row displays a scene-level 3D open-vocabulary map along with the grounding of ”parked trucks” on the map.
@@ V CONCLUSION
We propose LMPOcc, a novel 3D occupancy prediction framework that leverages long-term memory priors from historical traversals to enhance local perception and build unified global occupancy maps. By introducing a model-agnostic prior format, LMPOcc ensures compatibility across different prediction baselines, while a lightweight Current-Prior Fusion module adaptively integrates prior and current features. Validated on the Occ3D-nuScenes benchmark, LMPOcc achieves state-of-the-art local occupancy prediction, especially for static semantic classes. Additionally, it supports large-scale global occupancy construction via multi-vehicle crowdsourcing and facilitates 3D open-vocabulary map building through occupancy-derived dense depth. Future research could conduct in-depth investigations into dynamic object processing within prior information and develop learnable optimization frameworks for global occupancy.